FEL Upgrade Project Weekly Brief September 23-27, 2002

 

 

Highlights:

Commissioning has started!

We are approaching a major milestone in the FEL Upgrade Project. Monday, September 30, is the end date for both the 10 kW IR Upgrade construction project funded by ONR and the first phase of the 1 kW UV Upgrade project funded by AFRL. We are continuing a transition from Upgrade construction activities to commissioning activities that actually began in July. Discrete goals which the project team had for Sept. 30 included: (1) linac under vacuum; (2) upgraded injector RF system (100 kW) delivering high power RF; and (3) upgraded high voltage power supply for the gun(HVPS-600 kV @10 mA) delivering high voltage. As of this writing, the more than 80% of the linac is under vacuum with the two linac cryomodules and injector cryounit cooled to 2k and ready for start of final SRF commissioning. As of midnight last night, the Personnel Safety System (PSS) was recertified which put all power systems back under configuration control and allows commissioning activities to proceed. The injector RF system has already delivered more than 40 kW to waveguides shorted at the cryounit input. The interlocks and control power of the gun HVPS were successfully tested yesterday and today we achieved 100kV.

Yesterday, the drive laser for the gun was successfully recommissioned and delivered power above specifications. High priority items for continuation of commissioning activities includes final assembly and pump-down of the beam transport hardware between the injector cryounit and first linac cryomodule and gun/light box assembly.

Management:

The project monthly reports for August were completed this week along with estimates of the financial close-out of both the 10 kW Upgrade project the first phase of the UV Upgrade project, since funding authority for both of these projects ends next week on Sept. 30th. Both projects will close with no cost overruns. Schedule slippage on late delivery of sextupoles/octupoles remains and as discussed at the May 2002 project review, a work around is in progress so these items do not significantly impact the commissioning schedule. The first phase of the UV FEL project has delivered its scope on cost and schedule.

Accel. Div. ES&H and FEL Project Management signed off on a decision memorandum to extend FEL operations to 30 kW of optical output within the context of the present safety envelope for the CEBAF site Environmental Assessment. Copies were sent to the DOE Site Office.

 

FEL Readiness Process (FRP) for the 10 kW Upgrade is focusing on three EH&S systems: Personnel, Laser, and RadCon. Per the previous FEL Readiness Plan of 1998, other aspects of the Upgrade are being assessed for the Readiness Process by Jefferson Lab and the DOE Site Office. Findings of the FRP are addressed according to Jefferson Lab readiness policy: "Readiness Issues" must be resolved before proceeding to next phase and "Readiness Concerns" are added to maintenance schedules. The FRP and results will be presented to the FEL Review Committee in November.

WBS 4 (Injector):

Gun Assembly:

We developed a problem with the gun assembly this week and have incorporated a work around with minimal impact on the commissioning schedule. During the assembly of the upgrade gun cathode structure early this week, we noted an exfoliation of the high voltage conditioning film that is deposited onto the cathode ball. Previously, no deterioration has been observed with this film on stainless steel substrates including on a large (100cm2) test sample that was coated 3 years ago and stored in ambient conditions. After some analysis of our records and procedures, it was determined that the localized coating failure was due to the use of an Al support rod which supported the ball during deposition. The film adhesion on the opposite side of the ball was unaffected. The action plan that was initiated immediately after the noted failure was the following:

(1) the film was removed from the cathode ball using a stainless steel etch that was previously tested when we had to remove trial films from the cathode support tube in July and August.

(2) we began re-polishing the ball in preparation for a re-deposition

(3) the deposition chamber was readied for use including machining of a stainless steel support in place of the original aluminum support.

It is estimated that the ball preparation procedure will take a week followed by another week for the 2-step deposition process. The process chamber has recently been upgraded with a larger vacuum pump and we note that the geometry of the ball allows this deposition to be done relatively quickly compared to the extended cylindrical geometry of the cathode support tube. This two week delay impacts the final assembly and pump down of the gun, however the delay should not significantly impact the overall commissioning schedule.

Other injector progress:

The drive laser transport line was completed up to the last mirror can and pumped down this week. Installation of the final tube with the exit Brewster window must wait for light box installation. We managed to mount the ion pump on the light box and install the first of four mirrors in the box.

 

The instrument flange was completed by the shop. Ion pumps have been mounted onto the gun chamber. The vacuum line, which connects the bake pump to the gun was fabricated. The solenoid has had its field measurements completed and is ready for installation. The modified mirror holders for the light box were fabricated and the ion pump was mounted onto the system. The support tube and high and ground end corona rings were mounted into the ceramic stack.

The implanted coating on the ball cathode was found to be peeling off. This might have been caused by aluminum contamination from the support mount used in the implant chamber. The coating has been removed with chemicals, and the ball cathode is being re-polished. A stainless steel mount was fabricated to replace the aluminum mount in the next ball cathode implant process.

Gun HVPS - The HVPS for the Gun is assembled and passed PSS Certification. It was successfully operated at 100 kV in air today. An additional SF6 bag will be purchased to contain the additional volume of the new tank.

 

WBS 6 (RF):

Quarter HVPS - Unit 3 was turned On under local control and ran 44 kW of RF into a short at reduced cathode voltage. The voltage taps were then set to 95% of full voltage and achieved 34 kV with the klystron load of 6.4 amps. This voltage is more than enough to achieve 110 kW for the klystrons. The ripple voltage is 510 volts p-p at full power. Unit 3 also passed the PSS Certification. Unit 4 HVPS and klystron cart are still being assembled and will be tested within the next 2 weeks. The system prints are being updated and nearly ready for release.

Buncher - The Buncher water skid controller is not connected to the Buncher cavity or the LCW. This is to relieve congestion for the installation of the Gun and Light Box. It passed the PSS Certification with a jumper for the LCW interlock on the water skid.

Zone 2 - The waveguides for this zone were reconnected again this week. It passed the PSS Certification.

Zone 4 - A designer will determine, layout, and have the Machine Shop build the remaining waveguides for this zone. Zone 2 will be used as an example, but since the penetrations are not identical, the lengths will be slightly different. Due to a bad connection on the waveguide interlock, this zone did not pass the PSS Certification. The problem was fixed today.

WBS 8 (Instrumentation):

The interlocks to the gun high voltage power supply are now complete. The safety group did the PSS certification this week and we have begun to turn on the various systems. The level of activity has gone off the charts and these reports often do not reflect the accomplishments nor the superb efforts of the team. AC power to FL01B08 and FL05B07 and 08 has been completed with the breaker boxes upgraded to support 208, 3 phase power. These racks support the Injection and Extraction dipoles. The Injection and Extraction 5 kW Danfysiks power supplies also were installed in Zones 1 and 5 respectively. Water hookups to the PS have been completed. Shunt chassis for the First Arc String have been installed in FL05B09 and the wiring is in progress. Radcon has begun work in the vault for the Rapid Access system. Four GP 64 bit VME Isolation modules are being populated. Assembly has begun in the 0f and 1F regions for the Beamviewer camera assemblies.

BLMs (Beam Loss Monitors)- Testing of the prototypes is almost complete. The cards are being tested with the actual BLM heads and the LeCroy high-voltage power supply. They will next be calibrated and then installed. This will provide the baseline functionality for commissioning for 12 channels, after which a second set will be built-up and further developed to provide the advanced features such as the "beam veto during the macropulse" and the "trip sequencing" functions that are not fully developed. Once the firmware for these function has been developed, the 1st set will be 'flashed' with the upgrade.

Parts have been ordered for 40 additional beamviewer control boxes. And we are in the process of stuffing more of the GPIM64 VME boards for installation in various systems. EH&S staff and Deborah did a safety walkthrough, no major deficiencies were noted, however general housekeeping could use some improvement.

This week we began a collaboration with MIT Bates lab to get EPICS working on Linux. This has been done by many groups using the X86 chips but our focus will be with the new Motorola Power PC chip MPC8266ADS-PCI. In fact our group is using a Dell PC and framegrabber as

an IOC to process the viewer video. Two students whom have been writing EPICS code for 1.5 years came down for three days. We worked together to establish a clear path that will result in a low cost - license free IOC. This PPC chip has PCI bus support allowing the use of countless commercial I/O boards. We are setting up a website and mail list that is open to all interested: http://laser.jlab.org/linuxioc To subscribe to the mail list send email to linuxioc@jlab.org and put SUBSCRIBE in subject line.

WBS 9 (Transport):

Dipoles

Optical Chicane Dipoles (GW)

MEI is still getting the core pieces reground for the remaining 6 dipoles and is poised to do the assembly as soon as they are ready.

Injector Dipoles (DU/DV)

The first GV qualified for the machine and was signed off by Dave Douglas after two additional iterative magnet measurement learning cycles.

We had solved the linearity of the field integral as you measure in transverse positions (integral gradient) last week. We homed in on the values of that field integral in the transverse positions, obtaining the correct values to within 2 parts in 10,000. In doing so, we calibrated our spread sheet calculations to dial the micrometric field clamp adjustment values.

Formal measurements to this point explored and finally obtained the nominal 10 MeV/c requirements of 555.940 Gauss and a half sagitta chord integral of 11673.9 g-cm. When we went to explore the 9.5 MeV/c bulk field setting the field integral didn't scale by 1.8%- way out of Dave's specs. The dynamic range of these magnets is very small because they depend on hysteresis in the yoke. As a result, David Douglas redefined the specification for this magnet style to operate at 10,750 g-cm, (9.2 MeV/c) the value the IR Demo successfully operated at. We quickly homed in on this integral and the corresponding bulk field (by shifting the magnetically defined centerline of this wedge magnet) and obtained bulk field flatness and integral gradient to within Dave's requirements and signed off the magnet as qualified.

Arc 180 Degree Dipoles (GY)

GY Coil Potting fixture continues its preparation for potting.

Wang NMR continued winding the remaining two layers of the last GY coil.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

At Wang NMR, they started assembly of the GQ magnet

The first two GX coils were impregnated and cured

I will visit them next week to be there during the final assembly of the GQ.

Quadrupoles

Trim Quad (QT)

The third and forth trim quads were fully measured

Sextupole (SF)

We passed a preliminary layout of the quick fix sextupoles to Process equipment, the folks who had made the original IR Demo sextupole cores for comment on the shimming.

Octupole (OT)

The Octupole design continued at DULY Research where they generated the core and pole tip drawings.

Beam Line and Vacuum

X-Chamber - Fabrication of the chamber is finished and it is being cleaned and leak tested by the Vacuum Group.

Chicane Chambers - Work on other chamber plates, tubes, and CFF's are on hold to allow higher priority parts to be fabricated.

Y Chamber - Plate machining is in progress and is almost complete.

ARC Chambers - Master Machine's work on the first ARC chamber is nearly complete. They now have a good handle on their production rate (too slow) which could be increased substantially if with more welding capability. They are therefore working on getting a qualified second shift welder.

We signed off the location drawing for the extraction stands.

Girder and vacuum assembly continued

Low Conductivity Water piping installation continued in the FEL enclosure, it is now more than 60% complete.

WBS 11 (Optics):

The tungsten disulfide sputtered leadscrews for the optical cavities were tested under load this week. Initial tests looked quite good, with a measurement of drag torque of ~ 20 in-lbs. However, after a few iterations, the drag torque increased. Our initial impression, based on microscopic inspection, suggests the film is wearing off. While we will continue to investigate this, for now, we intend to move forward with the use of a Vespel nut/bare stainless steel leadscrew combination. The LVDTs have been checked for outgassing, and are sufficiently good for the IR optical cavity and transport system. Outgassing checks on other components are ongoing at the time of this report. We received the HR mirrors for 6 micron operations. They are being inspected. A prototype outcoupler mirror holder was nickel-plated prior to brazing of the cooling tube. Some chipping of an edge (in a noncritical area) was noted, and suggests we want to make that edge thicker. The OCMMS beamline hardware was sent to the machine shop for modifications to one item, and final UHV cleaning. When they return, they will be installed in the beamline. Procurement of optical transport hardware is proceeding at a rapid rate.

Commissioning activities:

In preparation for commissioning activities, a leak in the drive laser beam dump was repaired and the drive laser operated. It is operating at above spec power. Other specs will be checked next week. The drive laser optical transport was aligned from the drive laser enclosure to the last turning mirror above the light box. At this time the transport is being pumped out to eliminate water, and will be backfilled with dry nitrogen.